Grade 12th passMechanics

a block of mass m1 = 3.70 kg on a frictionles inclined plane of angle theta = 28.0 is connected by a cord over a small frictionless, massless pulley to a second block of mass m2 = 1.86 kg hanging vertically.a) what is the acceleration of each block .b)find the tension in the cord .

Arshia Shahid
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
 
Conceptually, you know that the blocks will have the SAME acceleration and the SAME tension.

This is easier to see with a Free-Body Diagram, but I’ll try to explain it as best as I can. You need to first pick a direction that the blocks are moving. For this part, it doesn’t matter which way you designate the motion, it is just for keeping our positive a negative labels consistent. I am going to state that the hanging block is moving DOWN and the other block is moving UP THE INCLINE.  

Now, just solve the following equation for both blocks: 

ΣF = m(a) 

Where: 
ΣF = sum of forces acting along the direction of motion 
m = mass 
a = acceleration 

BLOCK 1 
ΣF1 = m1(a) 

Where: 
ΣF1 = T – W1sin30 
m1 = mass of block 1 = 3.70kg 
a = acceleration = unknown 
T = tension = unknown 
W1 = weight of block 1 = m1(g) = 37 N 

T – W1(sin28) = m1(a) 
T – (37 N)(sin 28) = (3.70kg)a 
T – 17.37N = (3.70kg)(a) 

BLOCK 2 
ΣF2 = m2(a) 

Where: 
ΣF2 = W2 – T 
m2 = mass of block 2 = 1.86kg 
a = acceleration = unknown 
T = tension = unknown 
W2 = weight of block 2 = m2(g) = 18.6N 

W2 – T = m2(a) 
18.6N – T = (18.6kg)(a) 

Now, solve for T in the block 1 equation and plug that into the equation for block 2 

T = 17.37N + (3.70kg)(a) 

18.6N – (17.37N + (3.70kg)(a)) = (1.86Kg)(a) 
 
Now solve for a